# Certifiable Deep Learning for Reachability Using a New Lipschitz Continuous Value Function

### v2

@@ Abstract
We propose a new reachability learning framework for high-dimensional nonlinear systems, focusing on reach-avoid problems. These problems require computing the reach-avoid set, which ensures that all its elements can safely reach a target set despite any disturbance within pre-specified bounds. Our framework has two main parts: offline learning of a newly designed reach-avoid value function and post-learning certification. Compared to prior works, our new value function is Lipschitz continuous and its associated Bellman operator is a contraction mapping, both of which improve the learning performance. To ensure deterministic guarantees of our learned reach-avoid set, we introduce two efficient post-learning certification methods. Both methods can be used online for real-time local certification or offline for comprehensive certification. We validate our framework in a 12-dimensional crazyflie drone racing hardware experiment and a simulated 10-dimensional highway takeover example.
@@
Reachability, nonlinear control, deep reinforcement learning.
@@ I Introduction
Ensuring the safe and reliable operation of robotic systems in uncertain environments is a critical challenge. For instance, humanoid robots must safely navigate around humans while performing their tasks, and air taxi vehicles need real-time synthesis of safe, collision-free trajectories in crowded airspace. In these safety-critical applications, it is essential to complete various tasks safely despite unpredictable disturbances. Additionally, knowing the set of states from which the task can be safely completed is important because the task can fail once the robot exits that set. Reachability analysis addresses this challenge by computing the optimal control and determining the reach-avoid set—a set of states that can safely reach a target set under all possible disturbances within a specified bound.
Traditional Hamilton-Jacobi reachability analysis methods, such as [2, 3, 4], leverage dynamic programming to synthesize the optimal control and a reachability value function, whose sign indicates whether or not a state can safely reach the target set. Though theoretically sound, they suffer from the curse of dimensionality [5]. As the system’s dimension increases, the computational complexity grows exponentially, making these methods impractical for real-world applications without approximation or further logic to manage the problem size.
There has been interest in leveraging machine learning techniques to learn reachability value functions for high-dimensional systems [6, 7, 8, 9, 10, 11, 12]. However, a major drawback of existing learning-based approaches is the lack of deterministic guarantees. Without such guarantees, it is difficult to ensure the reliability and safety of the robotic systems in critical scenarios.
Recent works [13, 14] provide probabilistic safety guarantees for the learned reach-avoid sets. Additionally, safety filter approaches [11, 12, 15] have been proposed, which offer point-wise guarantees by ensuring safety for individual states.
In this work, we asked if it is possible to learn a set of states that have deterministic reach-avoid guarantees. We propose a new framework for learning trustworthy reach-avoid sets for high-dimensional systems. Our framework involves first learning a newly designed reach-avoid value function, and then conducting set-based certification to ensure that all states in a certified reach-avoid set safely reach the target set despite any disturbance. Specifically:
@@
  1. 1.
    We propose a new reach-avoid value function that is provably Lipschitz continuous, and its Bellman operator is a contraction mapping, removing the need for contractive Bellman operator approximation commonly used in prior works. Moreover, we demonstrate that the optimal control derived from our value function tends to reach the target set more rapidly than those from prior works. We apply deep reinforcement learning (RL) to learn this new value function for high-dimensional nonlinear systems.
  2. 2.
    We develop two reach-avoid set certification methods. The first uses the Lipschitz constant of the dynamics to certify the safety of a subset of the learned reach-avoid set, ensuring all its elements can safely reach the target set amid potential disturbances. The second employs second-order cone programming to do the same. Both methods offer deterministic assurances. They can be applied online to verify if a neighboring set around the current state can safely reach the target set despite any disturbance within a pre-specified set, or offline for comprehensive certification.
  3. 3.
    We demonstrate the computational benefits of our new value function and the assurance of our (real-time) certification methods through hardware experiments and simulations. We show empirically that our learned policy compares favorably with that of prior works, and that the Lipschitz continuity of our new value function can accelerate value function learning.
@@ II Related works
Reachability learning methods. DeepReach [6] is a pioneering work on learning finite-horizon reachability value functions. In other studies, such as [7, 8, 10, 9, 11, 15, 12], the assumption of a known horizon is relaxed and infinite horizon reachability learning problems are considered. In this work, we also consider the infinite horizon case. We introduce a new value function which is provably Lipschitz continuous and whose Bellman operator is a contraction mapping, offering favorable computational efficiency when compared with prior work.
Verification of learning-based control. Recent work [13, 14, 17] has provided probabilistic safety guarantees for DeepReach; in this paper, we introduce methods that provide deterministic reach-avoid guarantees and we show how they could be used locally in real time. Other studies [11, 15, 12] provide point-wise safety filters. However, we propose set-based reach-avoid certification methods to verify if all states in a set can safely reach the target set amid potential disturbances. Our certification methods also differ from existing set-based approaches for verifying neural network-controlled systems, including those that verify regions of attraction [18, 19, 20], forward reachability sets [21, 22, 23, 24], and safe sets using barrier certificates [25, 26, 27]. In general, these methods do not certify if a set of states is within the ground truth reach-avoid set, as ours do.
Constrained optimal control. Control barrier functions (CBFs) offer safety guarantees [28, 29, 30], but they tend to be more conservative than model predictive control (MPC) [31, 32], which optimally balances task performance and safety. However, constructing or learning CBFs and solving MPC can be difficult for nonlinear systems with complex constraints. By leveraging deep neural networks, constrained reinforcement learning (CRL) [33, 34, 35] learns control policies that maximize task rewards while adhering to complex constraints. Nonetheless, the positive sign of a CBF provides safety information for a state; it does not indicate whether a trajectory from that state can reach the target set. Similarly, the value functions in MPC and CRL do not provide information about whether a state can safely reach the target set. In contrast, our new value function not only provides the optimal control for the worst-case disturbance, but also indicates, based on its sign, whether or not a state is in the reach-avoid set.
@@ III Problem Formulation
We consider the uncertain nonlinear dynamics described by
$x_{t+1}=f(x_{t},u_{t},d_{t}),$ (1)
where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in{\mathcal{U}}\subseteq{\mathbb{R}^{m_{u}}}$ is the control, and $d_{t}\in{\mathcal{D}}\subseteq{\mathbb{R}^{m_{d}}}$ represents the disturbance, such as model mismatches or uncertain actions of other agents. We assume both ${\mathcal{U}}$ and ${\mathcal{D}}$ are compact and continuous sets. We define a state trajectory originating from $x_{0}$ under a control policy $\pi(\cdot):\mathbb{R}^{n}\to{\mathcal{U}}$ and a disturbance policy $\phi(\cdot,\cdot):\mathbb{R}^{n}\times{\mathcal{U}}\to{\mathcal{D}}$ as ${\xi_{x_{0}}^{\pi,\phi}}:=\{x_{t}\}_{t=0}^{\infty}$, where $x_{t+1}=f(x_{t},{\pi}(x_{t}),{\phi}(x_{t},{\pi}(x_{t})))$$\forall t$.
Let ${\mathcal{T}}\subseteq\mathbb{R}^{n}$ be an open set, representing a desired target set. We assume that there exists a Lipschitz continuous, bounded reward function $r(x)$ indicating if a state $x$ is in the target set,
$r(x)>0\Longleftrightarrow x\in{\mathcal{T}}.$ (2)
We consider a finite number of Lipschitz continuous, bounded constraint functions $c_{i}(x)>0,\forall i\in{\mathcal{I}}$. We can simplify the representation of constraints by considering their pointwise minimum $c(x):=\min_{i\in{\mathcal{I}}}c_{i}(x)$. We define the constraint set as ${\mathcal{C}}:=\{x:c(x)>0\}$, and we have
$c(x)>0\Longleftrightarrow x\in{\mathcal{C}}.$ (3)
@@
which includes all the states that can reach the target set safely in finite time despite any disturbance within the set ${\mathcal{D}}$.
In this paper, we compute the RA set for high-dimensional nonlinear systems with complex constraints.
@@
We consider a 12-dimensional dynamics [38], where the $i$-th drone’s state is $x_{t}^{i}=[p_{x,t}^{i},v_{x,t}^{i},p_{y,t}^{i},v_{y,t}^{i},p_{z,t}^{i},v_{z,t}^{i}]$. In each of the $(x,y,z)$ axes, the $i$-th drone is modeled by double integrator dynamics, and the control is its acceleration $u_{t}^{i}=[a_{x,t}^{i},a_{y,t}^{i},a_{z,t}^{i}]$, with $\|u_{t}^{i}\|_{\infty}\leq\varepsilon_{u}:=1\textrm{m/s}^{2}$. We model the center of the gate as the origin. The radius of the orange gate is 0.3 meters, and the radius of the crazyflie drone is 0.05 meters. We consider a target set for the ego drone, as plotted in Figure 1:
@@
IV A New RA Value Function
In this section, we propose a new RA value function for evaluating if a state belongs to the RA set. Unlike prior works [7, 8, 10, 9, 11, 15, 12], our value function incorporates a time-discount factor. This results in a Lipschitz-continuous value function, which appears to accelerate the learning process, and establishes a contractive Bellman equation, eliminating the need for the contractive Bellman equation approximation commonly used in prior works. Furthermore, we show that the optimal control derived from this new value function tends to reach the target set rapidly.
@@ IV A New RA Value Function
${\bar{V}}(x):=\max_{{\pi}}\min_{{\phi}}\sup_{t=0,\dots}{g(\xi_{x_{0}}^{\pi,\phi},t)}$ (8)
where ${\bar{V}}(x)>0$ if and only if $x\in{\mathcal{R}}$. We compute ${\bar{V}}(x)$ by solving its Bellman equation. However, ${\bar{V}}(x)$ has a non-contractive Bellman equation, whose solution may not recover ${\mathcal{R}}$, as shown in Figure 2. To address this, prior works include a time-discount factor $\gamma$ to create a contractive Bellman equation approximation. For each $\gamma\in(0,1)$, there is a unique solution to the approximated Bellman equation, which converges to ${\bar{V}}(x)$ as $\gamma$ is gradually annealed to 1.
Inspired by previous studies, we enhance computational efficiency by designing a new time-discounted RA value function. Ours incorporates a time-discount factor into the value function formulation, resulting in a contractive Bellman equation without the need for any approximation. This improvement eliminates the requirement of the $\gamma$-annealing process commonly used in prior works [7, 8, 10, 9, 11, 15, 12], and significantly reduces computation time.
@@
${g_{\gamma}}({\xi_{x_{0}}^{\pi,\phi}},t):=\min\Big{\{}\gamma^{t}r\big{(}{x_{t}}\big{)},\min_{\tau=0,\dots,t}\gamma^{\tau}c\big{(}{x_{\tau}}\big{)}\Big{\}}.$
@@
Let $\gamma\in(0,1)$ and $V:\mathbb{R}^{n}\to\mathbb{R}$ be an arbitrary bounded function. Consider the Bellman operator ${B_{\gamma}}[\cdot]$ defined as,
${B_{\gamma}}[V(x)]\coloneqq\max_{u}\min_{d}\min\big{\{}c(x),\max\{r(x),\gamma V(f(x,u,d))\}\big{\}}.$ (11)
Then, ${V_{\gamma}(x)}$ in (9) is the unique solution to the Bellman equation $V(x)={B_{\gamma}}[V(x)]$. We have $\|{B_{\gamma}}[{V_{\gamma}^{1}}]-{B_{\gamma}}[{V_{\gamma}^{2}}]\|_{\infty}\leq\gamma\|{V_{\gamma}^{1}}-{V_{\gamma}^{2}}\|_{\infty}$, for all bounded functions ${V_{\gamma}^{1}}$ and ${V_{\gamma}^{2}}$.
@@
The proof can be found in the technical report [1]. ∎
@@
The proof can be found in the technical report [1]. ∎
@@
Moreover, the optimal control derived from our value function reaches the target set quickly, as a trajectory that reaches the target set rapidly incurs a high time-discounted RA measure value.
@@
Let $x$ be in the RA set and ${\phi}$ be an arbitrary disturbance policy. Suppose $({\pi_{1}},{t_{1}})$ and $({\pi_{2}},{t_{2}})$ are two control policies and corresponding times to maximize the discounted RA measure $g_{\gamma}$, $\gamma\in(0,1]$,
$({\pi_{1}},{t_{1}}),({\pi_{2}},{t_{2}})\in\arg\max_{{\pi},t}g_{\gamma}(\xi_{x}^{{\pi},{\phi}},t).\vspace{-0.5em}$ (12)
If ${t_{1}}<{t_{2}}$, then for all $\check{\gamma}\in(0,\min\{\gamma,\frac{{V_{\gamma}(x)}}{\max_{x}r(x)}\})$, we have $g_{\check{\gamma}}(\xi_{x}^{{\pi_{2}},{\phi}},{t_{2}})<g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})$.
@@
From the definitions of $\check{\gamma}$, ${g_{\gamma}}$, and ${V_{\gamma}(x)}$, and the boundedness of $r(x)$, we have $g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{2}})\leq\check{\gamma}^{{t_{1}}}\check{\gamma}\max_{x}r(x)\leq\check{\gamma}^{{t_{1}}}V_{\gamma}(x)<\big{(}\frac{\check{\gamma}}{\gamma}\big{)}^{{t_{1}}}V_{\gamma}(x)\leq\big{(}\frac{\check{\gamma}}{\gamma}\big{)}^{{t_{1}}}g_{\gamma}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})\leq g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})$. This completes the proof. ∎
@@
While a small time-discount factor in ${V_{\gamma}(x)}$ offers numerous benefits, it is not conclusive that $\gamma$ should always be near zero. In theory, for all $\gamma\in(0,1)$, the super-zero level set of ${V_{\gamma}(x)}$ recovers the exact RA set. However, in practice, a near-zero $\gamma$ can lead to a conservatively estimated RA set, where a trajectory reaching the target set at a late stage may have a near-zero or even negative time-discounted RA measure due to numerical errors. We will explore the trade-offs of selecting various $\gamma$ values in Section VII-D.

### v5

# Certifiable Reachability Learning Using a New Lipschitz Continuous Value Function

@@ Abstract
We propose a new reachability learning framework for high-dimensional nonlinear systems, focusing on reach-avoid problems. These problems require computing the reach-avoid set, which ensures that all its elements can safely reach a target set despite disturbances within pre-specified bounds. Our framework has two main parts: offline learning of a newly designed reach-avoid value function, and post-learning certification. Compared to prior work, our new value function is Lipschitz continuous and its associated Bellman operator is a contraction mapping, both of which improve the learning performance. To ensure deterministic guarantees of our learned reach-avoid set, we introduce two efficient post-learning certification methods. Both methods can be used online for real-time local certification or offline for comprehensive certification. We validate our framework in a 12-dimensional crazyflie drone racing hardware experiment and a simulated 10-dimensional highway take-over example.
@@
Reachability, Machine Learning for Robot Control, Reinforcement Learning
@@ I Introduction
Ensuring the safe and reliable operation of robotic systems in uncertain environments is a critical challenge as autonomy is introduced into everyday systems. For instance, we would like humanoid robots to safely work close to humans. As a second example, new concepts for air taxis will need real-time synthesis of safe trajectories in crowded airspace. These safety-critical applications are typically characterized by sequences of tasks, and knowing the set of states from which the task can be safely completed despite unpredictable disturbances is important. Reachability analysis addresses this challenge by determining the reach-avoid set—a set of states that can safely reach a target set under all possible disturbances within a specified bound, as well as the corresponding control.
Traditional Hamilton-Jacobi reachability analysis methods [1, 2, 3] leverage dynamic programming to synthesize the optimal control and a reachability value function, whose sign indicates whether or not a state can safely reach the target set. Though theoretically sound, they suffer from the curse of dimensionality [4]: as the system’s dimension increases, the computational complexity grows exponentially, making these methods impractical for real-world applications without approximation or further logic to manage the problem size.
There has been interest in using machine learning techniques to estimate reachability value functions for high-dimensional systems [5, 6, 7, 8, 9, 10, 11]. However, a major drawback of existing reachability learning methods is the lack of deterministic safety guarantees. Recent work [12, 13] provides probabilistic safety guarantees for the learned reach-avoid sets. Additionally, safety filter approaches [10, 11, 14] have been proposed, which offer point-wise guarantees by ensuring safety for individual states.
In this work, we propose the first method for learning reach-avoid sets for high-dimensional nonlinear systems with deterministic assurances. Our method involves learning a new reach-avoid value function (Sections IV and V), and then conducting set-based certification to ensure that all states in the certified set safely reach the target set despite disturbances (Section VI). Specifically:
@@
1) We propose a new reach-avoid value function that is provably Lipschitz continuous. Though it is not based on a Lagrange-type objective function (cumulative rewards over time) as in classical reinforcement learning (RL) [16], we prove its Bellman equation is still a contraction mapping, removing the need for contractive Bellman equation approximation commonly used in prior reachability works [10, 11]. Moreover, the control policy derived from our value function tends to reach the target set rapidly. We apply deep RL to learn this new value function.
2) We develop two reach-avoid set certification methods. The first uses the Lipschitz constant of the dynamics to certify the safety of a subset of the learned reach-avoid set, ensuring all its elements can safely reach the target set under disturbances. The second employs second-order cone programming to do the same. Both methods offer deterministic assurances. They can be applied online to verify if a neighboring set around the current state can safely reach the target set, or offline for verifying the same for a larger user-defined set.
3) We show the computational benefits of our new value function and the assurance of our (real-time) certification methods through simulations and hardware experiments. We empirically justify that the Lipschitz continuity of our new value function can accelerate value function learning.
@@ II Related works
Hamilton-Jacobi reachability learning. DeepReach [5] is a pioneering work on learning finite-horizon reachability value functions. In other studies, such as [6, 7, 9, 8, 10, 14, 11], the assumption of a known horizon is relaxed and infinite horizon reachability learning problems are considered. In this work, we also consider the infinite horizon case. We introduce a new value function which is provably Lipschitz continuous and whose Bellman operator is a contraction mapping, offering computational efficiency when compared with prior work.
Verification of learning-based control. Recent work [12, 13, 17] has provided probabilistic safety guarantees for DeepReach. In this paper, we introduce methods that provide deterministic reach-avoid guarantees and we show how they could be used locally in real time. Other studies [10, 14, 11] provide point-wise safety filters. However, we propose set-based reach-avoid certification methods to verify if all states in a set can safely reach the target set under potential disturbances. Our certification methods also differ from existing set-based approaches for verifying neural network-controlled systems, including those that verify regions of attraction [18, 19, 20], forward reachability sets [21, 22, 23, 24, 25], and safe sets using barrier certificates [26, 27, 28]. To the best of the authors’ knowledge, our work is the first to certify if a set of initial states is within the ground truth reach-avoid set.
Constrained optimal control. Control barrier functions (CBFs) offer safety guarantees [29, 30, 31], but they tend to be more conservative than model predictive control (MPC) [32, 33], which optimally balances task performance and safety. However, constructing or learning CBFs and solving MPC can be difficult for nonlinear systems with complex constraints. By leveraging deep neural networks, constrained reinforcement learning (CRL) [34, 35, 36] learns control policies that maximize task rewards while adhering to complex constraints. CBFs, along with the typical value functions defined in MPC and CRL, do not provide the information about whether a state can safely reach the target set. In contrast, our new value function not only provides the optimal control for the worst-case disturbance, but also indicates, based on its sign, whether or not a state can safely reach the target set.
@@ III Problem Formulation
We consider uncertain nonlinear dynamics described by
$x_{t+1}=f(x_{t},u_{t},d_{t}),\vspace{-0.3em}$ (1)
where $x_{t}\in\mathbb{R}^{n}$ is the state, $u_{t}\in{\mathcal{U}}\subseteq{\mathbb{R}^{m_{u}}}$ is the control, and $d_{t}\in{\mathcal{D}}\subseteq{\mathbb{R}^{m_{d}}}$ represents the disturbance, such as model mismatch or uncertain actions of other agents. We assume that both ${\mathcal{U}}$ and ${\mathcal{D}}$ are compact and connected sets. The disturbance bound could be estimated from prior data or a physical model. We define a state trajectory originating from an initial condition $x_{0}$ under a control policy ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi}:\mathbb{R}^{n}\to{\mathcal{U}}$ and a disturbance policy ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi}:\mathbb{R}^{n}\times{\mathcal{U}}\to{\mathcal{D}}$ as ${\xi_{x_{0}}^{\pi,\phi}}:=\{x_{t}\}_{t=0}^{\infty}$, where $x_{t+1}=f(x_{t},{\pi}(x_{t}),{\phi}(x_{t},{\pi}(x_{t})))$, $\forall t\in\{0,1,2,\cdots\}$.
Let ${\mathcal{T}}\subseteq\mathbb{R}^{n}$ be an open set, representing a target set. We assume that there exists a Lipschitz continuous, bounded reward function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}:\mathbb{R}^{n}\to\mathbb{R}$ indicating if a state $x$ is in the target set,
$r(x)>0\Longleftrightarrow x\in{\mathcal{T}}.\vspace{-0.3em}$ (2)
We consider a finite number of Lipschitz continuous, bounded constraint functions $c_{i}(x)>0,\forall i\in{\mathcal{I}}$, where ${\mathcal{I}}$ is the set of indexes of constraints. Throughout this paper, we define Lipschitz continuity with respect to the $\ell_{2}$ norm. We can simplify the representation of constraints by considering their pointwise minimum $c(x):=\min_{i\in{\mathcal{I}}}c_{i}(x)$. We define the constraint set as ${\mathcal{C}}:=\{x\in\mathbb{R}^{n}:c(x)>0\}$, and we have
$c(x)>0\Longleftrightarrow x\in{\mathcal{C}}.\vspace{-0.3em}$ (3)
@@
which includes all the states that can reach the target set safely in finite time despite disturbances within the set ${\mathcal{D}}$.
@@
We consider a 12-dimensional dynamics [39], where the $i$-th drone’s state is $x_{t}^{i}=[p_{x,t}^{i},v_{x,t}^{i},p_{y,t}^{i},v_{y,t}^{i},p_{z,t}^{i},v_{z,t}^{i}]$. In each of the $(x,y,z)$ axes, the $i$-th drone is modeled by double integrator dynamics, and the control is its acceleration $u_{t}^{i}=[a_{x,t}^{i},a_{y,t}^{i},a_{z,t}^{i}]$, with $\|u_{t}^{i}\|_{\infty}\leq\varepsilon_{u}:=1\textrm{m/s}^{2}$. We model the center of the gate as the origin. The radius of the orange gate is 0.3 meters, and the radius of the crazyflie drone is 0.05 meters. We consider a target set for the ego drone:
@@
IV A New Reach-Avoid Value Function
In this section, we propose a new RA value function for evaluating if a state belongs to the RA set. Unlike prior works [6, 7, 9, 8, 10, 14, 11], our value function incorporates a time-discount factor. This results in a Lipschitz-continuous value function, which appears to accelerate the learning process, and establishes a contractive Bellman equation, eliminating the need for the contractive Bellman equation approximation commonly used in prior works [6, 7, 9, 8, 10, 14, 11]. Furthermore, we show that the control policy derived from this new value function tends to reach the target set rapidly.
@@ IV A New Reach-Avoid Value Function
${\bar{V}}(x):=\max_{{\pi}}\min_{{\phi}}\sup_{t=0,\dots}{g(\xi_{x_{0}}^{\pi,\phi},t)},$ (9)
where ${\bar{V}}(x)>0$ if and only if $x\in{\mathcal{R}}$. We compute ${\bar{V}}(x)$ by solving its Bellman equation. However, ${\bar{V}}(x)$ has a non-contractive Bellman equation, whose solution may not recover ${\mathcal{R}}$, as shown in Figure 2. To address this, prior works [6, 7, 9, 8, 10, 14, 11] include a time-discount factor $\gamma$ to create a contractive Bellman equation approximation. For each $\gamma\in(0,1)$, there is a unique solution to the approximated Bellman equation, which converges to ${\bar{V}}(x)$ as $\gamma$ is gradually annealed to 1.
Inspired by previous studies, we enhance computational efficiency by designing a new time-discounted RA value function. Ours incorporates a time-discount factor into the value function formulation, resulting in a contractive Bellman equation without the need for any approximation. This improvement eliminates the requirement of the $\gamma$-annealing process commonly used in prior works [6, 7, 9, 8, 10, 14, 11], where $N$ approximated Bellman equations are solved sequentially with $\gamma$ values converging to 1. Theoretically, computing our new value function requires only $\frac{1}{N}$ of the time needed in prior works.
@@
${g_{\gamma}}({\xi_{x_{0}}^{\pi,\phi}},t):=\min\Big{\{}\gamma^{t}r\big{(}{x_{t}}\big{)},\min_{\tau=0,\dots,t}\gamma^{\tau}c\big{(}{x_{\tau}}\big{)}\Big{\}}.\vspace{-0.3em}$ (10)
@@
Let $\gamma\in(0,1)$ and $V:\mathbb{R}^{n}\to\mathbb{R}$ be an arbitrary bounded function. Consider the Bellman operator ${B_{\gamma}}[V]$ defined as,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{B_{\gamma}}[V](x)}\coloneqq\max_{u}\min_{d}\min\big{\{}c(x),\max\{r(x),\gamma V(f(x,u,d))\}\big{\}}.$
Then, we have $\|{B_{\gamma}}[{V_{\gamma}^{1}}]-{B_{\gamma}}[{V_{\gamma}^{2}}]\|_{\infty}\leq\gamma\|{V_{\gamma}^{1}}-{V_{\gamma}^{2}}\|_{\infty}$, for all bounded functions ${V_{\gamma}^{1}}$ and ${V_{\gamma}^{2}}$, and ${V_{\gamma}(x)}$ in (11) is the unique solution to the Bellman equation $V(x)={B_{\gamma}}[V](x)$.
@@
Let ${\pi}^{*}$ and ${\phi}^{*}$ be the optimal control and the worst-case disturbance policies. Observe ${V_{\gamma}}(x_{0})=\max_{\pi}\min_{\phi}\min\big{\{}c(x_{0}),\max\{r(x_{0}),\gamma\sup_{\tau=0,\dots}{g}(\xi_{x_{1}}^{{\pi}^{*},{\phi}^{*}},\tau)\}\big{\}}$ $=\max_{\pi}\min_{\phi}\min\big{\{}c(x_{0}),\max\{r(x_{0}),\gamma{V_{\gamma}}(x_{1})\}\big{\}}$, where $x_{1}=f(x_{0},\pi(x_{0}),\phi(x_{0},\pi(x_{0})))$ is the only variable affected by ${\pi}$ and ${\phi}$. Following [16, p.234], it can be rewritten as ${V_{\gamma}}(x_{0})=\max_{u_{0}}\min_{d_{0}}\min\big{\{}c(x_{0}),\max\{r(x_{0}),\gamma{V_{\gamma}}(x_{1})\}\big{\}}.$
Thus, ${V_{\gamma}}(x)$ is a valid solution to the Bellman equation $V={B_{\gamma}}[V]$. We show it is a unique solution by proving that ${B_{\gamma}}[V]$ is a contraction mapping when $\gamma\in(0,1)$, i.e., $\|{B_{\gamma}}[V_{1}]-{B_{\gamma}}[V_{2}]\|_{\infty}\leq\gamma\mathopen{}\mathclose{{}\left\|V_{1}-V_{2}}\right\|_{\infty}$, where $V_{1}$ and $V_{2}$ are two arbitrary bounded functions. Let $x$ be an arbitrary state. We have $\|{B_{\gamma}}[V_{1}](x)-{B_{\gamma}}[V_{2}](x)\|_{\infty}\leq\|\gamma\max_{u}\min_{d}V_{1}(f(x,u,d))-\gamma\max_{u}\min_{d}V_{2}(f(x,u,d))\|_{\infty}$. Since the max-min operator is non-expansive, we have, for all $x$, $\|{B_{\gamma}}[V_{1}](x)-{B_{\gamma}}[V_{2}](x)\|_{\infty}\leq\gamma\|V_{1}(x)-V_{2}(x)\|_{\infty}$. ∎
@@
Consider two arbitrary initial states $x_{0},x_{0}^{\prime}\in\mathbb{R}^{n}$. Let $\pi^{*}$ and $\phi^{*}$ be the optimal control and the worst-case disturbance policies. For each $t\in\{0,1,2,\dots\}$, define $x_{t+1}:=f(x_{t},u_{t},d_{t}^{\prime})$, $u_{t}:=\pi^{*}(x_{t})$, $x_{t+1}^{\prime}:=f(x_{t}^{\prime},u_{t},d_{t}^{\prime})$, and $d_{t}^{\prime}:=\phi^{*}(x_{t}^{\prime},u_{t})$. Let $\mathbf{u}:=\{u_{t}\}_{t=0}^{\infty}$, $\mathbf{d}^{\prime}:=\{d_{t}^{\prime}\}_{t=0}^{\infty}$, $\mathbf{x}:=\{x_{t}\}_{t=0}^{\infty}$, and $\mathbf{x}^{\prime}:=\{x_{t}^{\prime}\}_{t=0}^{\infty}$. Given an arbitrarily small $\varepsilon>0$, there exists a $\bar{t}<\infty$ such that $V_{\gamma}(x_{0})\leq g_{\gamma}(\mathbf{x},\bar{t})+\varepsilon.$ By definition, we have $V_{\gamma}(x_{0}^{\prime})\geq g_{\gamma}(\mathbf{x}^{\prime},\bar{t})$. Combining two inequalities, we have $V_{\gamma}(x_{0}^{\prime})-V_{\gamma}(x_{0})+\varepsilon\geq\min\{\gamma^{\bar{t}}(r(x^{\prime}_{\bar{t}})-r(x_{\bar{t}})),\min_{\tau=0,...,\bar{t}}\gamma^{\tau}(c(x^{\prime}_{\bar{t}})-c(x_{\bar{t}}))\}\geq-\max\{L_{r}\gamma^{\bar{t}}L_{f}^{\bar{t}},\max_{\tau=0,...,\bar{t}}L_{c}\gamma^{\tau}L_{f}^{\tau}\}\|x_{0}-x_{0}^{\prime}\|_{2}.$ The condition $\gamma L_{f}<1$ implies $(\gamma L_{f})^{\bar{t}}<1$, $\forall\bar{t}$. As a result, $V_{\gamma}(x^{\prime}_{0})-V_{\gamma}(x_{0})+\varepsilon\geq-L\|x_{0}-x_{0}^{\prime}\|_{2}$. Similarly, we can show that $V_{\gamma}(x_{0})-V_{\gamma}(x_{0}^{\prime})+\varepsilon\geq-L\|x_{0}-x_{0}^{\prime}\|_{2}.$ Combining these two inequalities, we prove Theorem 2. ∎
@@
Moreover, the control policy derived from ${V_{\gamma}(x)}$ reaches the target set quickly, as a trajectory that reaches the target set rapidly incurs a high time-discounted RA measure value.
@@
Let $x$ be in the RA set and ${\phi}$ be an arbitrary disturbance policy. Let $\gamma\in(0,1)$. Suppose $({\pi_{1}},{t_{1}})$ and $({\pi_{2}},{t_{2}})$ are two control policies and corresponding times to maximize the discounted RA measure $g_{\gamma}$, $({\pi_{1}},{t_{1}}),({\pi_{2}},{t_{2}})\in\arg\max_{{\pi},t}g_{\gamma}(\xi_{x}^{{\pi},{\phi}},t).$ Moreover, suppose ${t_{1}}<{t_{2}}$, then for all $\check{\gamma}\in(0,\min\{\gamma,\frac{{V_{\gamma}(x)}}{\max_{x}r(x)}\})$, we have $g_{\check{\gamma}}(\xi_{x}^{{\pi_{2}},{\phi}},{t_{2}})<g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})$.
@@
From the definitions of $\check{\gamma}$, ${g_{\gamma}}$, and ${V_{\gamma}(x)}$, and the boundedness of $r(x)$, we have $\check{\gamma}<1$, and $g_{\check{\gamma}}(\xi_{x}^{{\pi_{2}},{\phi}},{t_{2}})\leq\check{\gamma}^{{t_{1}}}\check{\gamma}\max_{x}r(x)\leq\check{\gamma}^{{t_{1}}}V_{\gamma}(x)<\big{(}\frac{\check{\gamma}}{\gamma}\big{)}^{{t_{1}}}V_{\gamma}(x)\leq\big{(}\frac{\check{\gamma}}{\gamma}\big{)}^{{t_{1}}}g_{\gamma}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})\leq g_{\check{\gamma}}(\xi_{x}^{{\pi_{1}},{\phi}},{t_{1}})$. ∎
@@
While a small time-discount factor in ${V_{\gamma}(x)}$ offers numerous benefits, it is not conclusive that $\gamma$ should always be near zero. In theory, for all $\gamma\in(0,1)$, the super-zero level set of ${V_{\gamma}(x)}$ recovers the exact RA set. However, in practice, a near-zero $\gamma$ can lead to a conservatively estimated RA set, where a trajectory reaching the target set at a late stage may have a near-zero or even negative time-discounted RA measure due to numerical errors. We will explore the trade-offs of selecting various $\gamma$ values in Section VIII-D.